# Asymgenesis

### v2

@@ Abstract
Table 1: The numerical coefficients $x_{C}$ (brown) and $y_{C}$ (green), associated with the conserved charges $\bar{\mu}_{C}=q_{C}6/T^{2}$ as defined in Eq. (6) for the scenario with $\mathcal{O}_{C}^{\rm{(SM)}}=\overline{L_{e}}He_{R}$ and $\mathcal{O}_{X}^{(D)}=X^{p}$ in different temperature regimes. The first column shows $T_{\rm{dec}}$, which is identified with $T_{B-L}$ $(T_{\rm{eq}})$ for $\bar{\mu}_{B-L}^{\rm eq}$ $\left(\mu_{X}^{\rm eq}\right)$, see text below Eq. (6), and (i)-(v) indicate the following ranges for $T_{\rm{dec}}$ [GeV]: $\quantity(10^{5},10^{6})$, $\quantity(10^{6},10^{9})$, $\quantity(10^{9},10^{11-12})$, $\quantity(10^{11-12},10^{13})$, $\quantity(10^{13},10^{15})$, respectively. The flavor index $\alpha$ in the second column lists the uncorrelated charged-lepton flavors that interact with $N_{1}$ in the relevant temperature range, assuming the SM–DM portal is active (decoupled) in that range. In the latter case, the results for $x_{C}$ reduce to those in standard WILG, c.f. Table II in Domcke:2020quw , which are included in parentheses in column 333 below. The ✗ symbol marks the absence of the corresponding $\bar{\mu}_{C}$ due to an efficient SM interaction. The coefficients $P$ and $P_{e,\tau}$ encode the primordial $q_{e,\mu,\tau}$ asymmetries’ flavor composition relative to the $N_{1}$ washout direction, as detailed in Domcke:2020quw . See text for further details.
$T_{\rm{dec}}$ Index $\alpha$
$\bar{\mu}_{Xe}$ ($\bar{\mu}_{e}$)
$\bar{\mu}_{2B_{1}-B_{2}-B_{3}}$
$\bar{\mu}_{u-d}$
$\bar{\mu}_{d-s}$
$\bar{\mu}_{B_{1}-B_{2}}$
$\bar{\mu}_{\mu}$
$\bar{\mu}_{u-c}$
$\bar{\mu}_{\tau}$
$\bar{\mu}_{d-b}$
$\bar{\mu}_{B}$
$\bar{\mu}_{u}$
$\bar{\mu}_{\Delta_{\perp}}$
$\bar{\mu}_{\Delta_{\alpha}}$
(i) $e,\mu,\tau$ $(e,\mu,\tau)$
$\frac{-3p^{2}}{22+10p^{2}}$ $(\frac{-3}{10})$,$\frac{-711}{1422+481p^{2}}$
[…]
(ii) $e,\mu,\tau$ $(e,\mu,\tau)$
$\frac{-3p^{2}}{38+17p^{2}}$ $\left(\frac{-3}{17}\right)$,$\frac{-1071}{2142+716p^{2}}$
0 (0)(0)( 0 ), 0
$\frac{-16-7p^{2}}{38+17p^{2}}$ $\left(\frac{-7}{17}\right)$, $\frac{45}{2142+716p^{2}}$
[…]
(iii) $e,\mu,\tau$ $(\parallel_{\tau},\tau)$
$\frac{-3p^{2}}{62+27p^{2}}$ $\left(\frac{142-225P_{\tau}}{247}\right)$,$\frac{-1779}{3558+1178p^{2}}$
0 (0)(0)( 0 ),0
$\frac{-6(7+3p^{2})}{62+27p^{2}}$ $\left(\frac{-123}{247}\right)$,$\frac{135}{3558+1178p^{2}}$
$\frac{-4(7+3p^{2})}{62+27p^{2}}$$\left(\frac{-82}{247}\right)$,$\frac{45}{1779+589p^{2}}$
$\frac{21+9p^{2}}{62+27p^{2}}$ $\left(\frac{123}{494}\right)$,$\frac{-135}{7116+2356p^{2}}$
$\frac{-3(2+p^{2})}{62+27p^{2}}$ $\left(\frac{142-225P_{\tau}}{247}\right)$,$\frac{-6}{1779+589p^{2}}$
$\left(\frac{225}{247}\right)$,
[…]
(iv) $e,\parallel_{e}$ ()($\parallel$)( ∥ )
$\frac{-9p^{2}}{102+37p^{2}}$ $\left(\frac{-23P+7}{30}\right)$,$\frac{-38}{76+23p^{2}}$
$\frac{12(3+p^{2})}{102+37p^{2}}$ $\left(\frac{1}{5}\right)$,$\frac{-3}{76+23p^{2}}$
$\frac{-36(3+p^{2})}{102+37p^{2}}$ $\left(\frac{-3}{5}\right)$,$\frac{9}{76+23p^{2}}$
$\frac{-10(3+p^{2})}{102+37p^{2}}$ $\left(\frac{-1}{6}\right)$,$\frac{5}{152+46p^{2}}$
$\frac{-18(3+p^{2})}{102+37p^{2}}$ $\left(\frac{-3}{10}\right)$,$\frac{9}{152+46p^{2}}$
$\frac{42+p^{2}(14-23P_{e})-60P_{e}}{102+37p^{2}}$ $\left(\frac{-23P+7}{30}\right)$,$\frac{-7}{152+46p^{2}}$
$\frac{18(3+p^{2})}{102+37p^{2}}$ $\left(\frac{3}{10}\right)$,$\frac{-9}{152+46p^{2}}$
$\frac{42+p^{2}(14-23P_{e})-60P_{e}}{102+37p^{2}}$ $\left(\frac{-23P+7}{30}\right)$,$\frac{-7}{152+46p^{2}}$
$\frac{-16(3+p^{2})}{102+37p^{2}}$ $\left(\frac{-4}{15}\right)$,$\frac{4}{76+23p^{2}}$
$\frac{120+46p^{2}}{306+111p^{2}}$ $\left(\frac{23}{90}\right)$,$\frac{23/6}{76+23p^{2}}$
$\frac{60+23p^{2}}{102+37p^{2}}$ $\left(\frac{23}{30}\right)$,$\frac{7}{152+46p^{2}}$
$\frac{[-16,7]}{152+46p^{2}}$
(v) $e,\parallel_{e}$ ()($\parallel$)( ∥ )
$\frac{-p^{2}}{14+5p^{2}}$ $\left(\frac{-3P+1}{4}\right)$,$\frac{-5}{10+3p^{2}}$
$\frac{4(3+p^{2})}{42+15p^{2}}$ $\left(\frac{1}{6}\right)$, $\frac{-1}{30+9p^{2}}$
$\frac{-20(3+p^{2})}{42+15p^{2}}$$\left(\frac{-5}{6}\right)$,$\frac{5}{30+9p^{2}}$
$\frac{-2(3+p^{2})}{14+5p^{2}}$ $\left(\frac{-1}{4}\right)$, $\frac{1}{20+6p^{2}}$
$\frac{-2(3+p^{2})}{14+5p^{2}}$ $\left(\frac{-1}{4}\right)$,$\frac{1}{20+6p^{2}}$
$\frac{6+p^{2}(2-3P_{e})-8P_{e}}{14+5p^{2}}$ $\left(\frac{-3P+1}{4}\right)$,$\frac{-1}{20+6p^{2}}$
$\frac{2(3+p^{2})}{14+5p^{2}}$ $\left(\frac{1}{4}\right)$,$\frac{-1}{20+6p^{2}}$
$\frac{6+p^{2}(2-3P_{e})-8P_{e}}{14+5p^{2}}$ $\left(\frac{-3P+1}{4}\right)$, $\frac{-1}{20+6p^{2}}$
$\frac{-8(3+p^{2})}{42+15p^{2}}$ $\left(\frac{-1}{3}\right)$,$\frac{2}{30+9p^{2}}$
$\frac{10+4p^{2}}{42+15p^{2}}$ $\left(\frac{1}{6}\right)$,$\frac{2}{30+9p^{2}}$
$\frac{8(3+p^{2})}{42+15p^{2}}$ $\left(\frac{1}{3}\right)$, $\frac{-2}{30+9p^{2}}$
$\frac{8+3p^{2}}{14+5p^{2}}$ $\left(\frac{3}{4}\right)$,$\frac{1}{20+6p^{2}}$
[…]
@@
Asymgenesis —  Consider a general scenario where primordial charges in both the visible and the dark sector can take arbitrary values and the charge-transfer interaction is in thermal equilibrium at high temperatures. The $B-L$ asymmetry in the strong wash-in regime and the ADM asymmetry can then, in full generality, be expressed as
@@
where $T_{B-L}$ $(T_{\rm{eq}})$ denotes the temperature when RHN interactions (the charge-transfer interaction) become(s) inefficient. Here, $\Delta_{\alpha}=B/3-L_{\alpha}$ denote lepton-flavor asymmetries corresponding to the flavor combinations that are decohered at the relevant temperature scale. At temperatures below $T_{\rm{eq}}$, both $q_{e}$ and $q_{X}$ are conserved separately. In particular, for $T_{B-L}<T_{\rm{eq}}$, $C=Xe$ can be replaced with $C=e$ in Eq. (6), with $q_{e}$ given by
@@
In this case, the numerical values for $x_{C}$ reduce to the result from standard WILG Domcke:2020quw . The values for $x_{C}$ and $y_{C}$ are shown for various temperature regimes in Table 1 for our example where $\mathcal{O}_{C}^{\rm{(SM)}}=\overline{L_{e}}He_{R}$ footnote . The coefficients without parentheses correspond to $T_{\rm{CG}}>T_{B-L}>T_{\rm{eq}}$, while the coefficients in parentheses correspond to $T_{\rm{CG}}>T_{\rm{eq}}>T_{B-L}$, where $T_{\rm{CG}}$ is the temperature scale associated with CG. The table shows that both the asymmetry in $X$ particles and the BAU are typically related to the primordial charge asymmetries generated during CG by an $\mathcal{O}(0.1-1)$ number. The analogous statement in the context of standard ADM models is that if the $B-L$ and $X$-violating higher-dimensional operator is in thermal equilibrium at high temperatures, then the BAU is related to the asymmetric component of DM by a $\mathcal{O}(1)$ number, barring potential washout processes. In conclusion, the framework considered so far predicts DM masses of $\mathcal{O}(1)-\mathcal{O}(10)$ GeV, which is similar to standard ADM models.
Beyond thermal equilibrium: SM to DM — Up to this point, we have analyzed the exchange of charges between the SM and the dark sector under the assumption that the charge-transfer interaction attained thermal equilibrium in the early universe. This requirement, however, is not essential. The same dynamics can be treated in an out-of-equilibrium framework using semiclassical Boltzmann equations. We will first consider the scenario where a CG mechanism generates primordial charges in the SM sector, and derive equations describing charge transfer to the DM sector. To facilitate this description, we introduce the following quantities
@@
The washout term, $W(z)$, associated with the interactions induced by Eq. (1) takes the following form,
@@
$Y_{\Delta X}(z)$ $=\frac{1}{\omega(z)}\left(Y_{\Delta X}(z_{\text{CG}})+\int_{z_{\text{CG}}}^{z}\omega(\tilde{z})W(\tilde{z})Y_{\Delta X}^{\rm{eq}}(\tilde{z})\,\mathrm{d}\tilde{z}\right),$ (12)
@@
$\omega(z)$ $=\exp\left(\int_{z_{\text{CG}}}^{z}W(\tilde{z})\,\mathrm{d}\tilde{z}\right),$ (13)
and $z_{\rm{CG}}=M_{N}/T_{\rm{CG}}$. As the electron-Yukawa interaction comes into equilibrium, $Y_{\Delta X}^{\rm{eq}}$ goes to zero. If, at that stage, the charge-transfer interaction characterized by $\Gamma^{w}$ remains efficient, Eqs. (12) and (13) show that $Y_{\Delta X}$ is driven to zero, thereby washing out the $X$ asymmetry. Meanwhile, if the charge-transfer interactions are initially in thermal equilibrium and decouple above the electron Yukawa equilibration temperature, $T_{e}\approx 85$ TeV Bodeker:2019ajh , then the final ADM yield, $Y_{\Delta X}^{\rm{end}}$, is driven exponentially close to $Y_{\Delta X}^{\rm{eq}}$, as given by Eq. (10) and Table 1. The attraction of $Y_{\Delta X}^{\rm{end}}$ towards $Y_{\Delta X}^{\rm{eq}}$ driven by $W$, can be referred to as dark wash-in, see also Asadi:2025vli .
Suppose instead that the effective-operator scale is so high that the charge-transfer interaction never enters thermal equilibrium. The charge-transfer then becomes a UV freeze-in (FI) process Giudice:2000ex ; Hall:2009bx ; Elahi:2014fsa . In such a case, $Y_{\Delta X}^{\rm{end}}$ is to a good approximation, determined by the value of $Y_{\Delta X}^{\rm{eq}}$ right after CG, and the value of $\Lambda$ and $T_{\rm{CG}}$. To unify the UV FI regime and the thermal-equilibrium regime, we find it useful to express the final asymmetry in $\Delta X$ as
@@
where $z^{\prime}=\text{Max}[z_{\rm{CG}},z_{\text{eq}}]$ and $\kappa$ parametrizes the charge-transfer efficiency. Numerical values for $\kappa$ for the case, $\mathcal{O}_{X}^{(D)}=X^{2}$, are shown in Fig. 1, where the horizontal and vertical axes show $T_{\rm{CG}}$ and $\Lambda$, respectively, and the black solid curve shows $H(z_{\rm{CG}})=\Gamma^{w}(z_{\rm{CG}})$. The latter marks the transition between the two regimes: The area below the curve corresponds to the thermal-equilibrium regime, $\Gamma^{w}(T_{\rm CG})/H(T_{\rm CG})>1$, where $\kappa\simeq 1$ and the asymmetry is well described by Tab. 1. The area above the curve shows the UV FI regime, $\Gamma^{w}(T_{\rm CG})/H(T_{\rm CG})<1$, and in the limit $\Gamma^{w}(T_{\rm CG})/H(T_{\rm CG})\ll 1$, we have $\kappa\simeq\Gamma^{w}(T_{\rm CG})/H(T_{\rm CG})$. For charge-transfer from the SM to the DM sector, the asymmetry in the dark sector is significantly smaller than the BAU in large portions of parameter space where Asymgenesis proceeds via UV FI. The black-dashed curve shows the contour corresponding to $\kappa=10^{-3}$, which would point to DM masses of $\mathcal{O}(1)$ TeV. This roughly represents an upper bound on ADM masses when the symmetric annihilation cross section is required to be large enough for the ADM component to be the dominant contribution to the DM relic density while respecting perturbative unitarity Griest:1989wd ; Cohen:2010kn ; Petraki:2013wwa . Interestingly, for $T_{\rm{CG}}=10^{14-15}$ GeV, it could be possible to account for the DM abundance even if the scale of NP associated with charge transfer is related to the scale of grand unification or possibly quantum-gravity effects, c.f. the upper-right corner of the figure. In summary, the region below the black solid curve in Fig. 1 corresponds to DM masses in the range of $\mathcal{O}(1)-\mathcal{O}(10)$ GeV, while the area below the black dashed curve corresponds to DM masses up to $\mathcal{O}(1)$ TeV. Above the black dashed curve, the symmetric component cannot annihilate efficiently due to the constraints imposed by perturbative unitarity.
Figure 1: Numerical result for $\kappa$ $(\tilde{\kappa})$, defined in Eq. (14) (Eq. (19)) for charge transfer from the SM (DM) sector to the DM (SM) sector, in the $T_{\rm{CG}}-\Lambda$ plane. See text for details.
@@
$Y_{\Delta e}(z)$ $=\frac{1}{\tilde{\omega}(z)}\left(Y_{\Delta e}(z_{\text{CG}})+\int_{z_{\text{CG}}}^{z}\tilde{\omega}(\tilde{z})\widetilde{W}(\tilde{z})Y_{\Delta e}^{\rm{eq}}(\tilde{z})\,\mathrm{d}\tilde{z}\right),$ (17)
@@
$\tilde{\omega}(z)$ $=\exp\left(\int_{z_{\text{CG}}}^{z}\widetilde{W}(\tilde{z})\,\mathrm{d}\tilde{z}\right).$ (18)
@@
In the scenario where the portal interaction attains thermal equilibrium, we have $Y_{\Delta e}^{\rm{eq}}(z^{\prime})\sim Y_{Xe}\sim Y_{\Delta X}(z^{\prime})$. This condition implies that, if the BAU is explained via WILG, the ADM mass must lie in the range $\mathcal{O}(1)-\mathcal{O}(10)$ GeV. In the FI regime, where $\tilde{\kappa}\ll 1$, the final ADM yield is typically much larger than the BAU, and one would expect ADM masses in the range $\mathcal{O}(\tilde{\kappa})-\mathcal{O}(10\,\tilde{\kappa})$ GeV. Specifically, the black dashed line in Fig. 1 corresponds to DM masses in the MeV range.
Finally, we note that since Asymgenesis builds on the type-I seesaw extension of the SM, it might be interesting to apply the framework to BSM models with gauged $U(1)_{B-L}$ symmetry. Contrary to conventional ADM models, Asymgenesis can accommodate scenarios where the charge-transfer mechanism decouples at temperatures that are orders of magnitude above the $B-L$-breaking scale, and still result in comparable abundances of baryonic matter and ADM.
Conclusion — In this work, we have proposed Asymgenesis, a framework based on the type-I seesaw model that can be used to account for the BAU and the observed DM density from initial conditions appropriate to either ADM or WILG. Compared to standard ADM scenarios that rely on higher-dimensional operators, Asymgenesis offers significantly greater flexibility in the charge-transfer interactions between the dark and visible sectors. Crucially, it disentangles the temperature at which the charge-transfer interactions decouple from the temperature scale at which $B-L$ violation becomes effective. By combining the advantages of WILG and ADM, the Asymgenesis framework considerably widens the model-building landscape for asymmetric dark matter.

### v3

@@ Abstract
Table 1: The numerical coefficients $x_{C}$ (brown) and $y_{C}$ (green), associated with the conserved charges $\bar{\mu}_{C}=q_{C}6/T^{2}$ as defined in Eq. (6) for the scenario with $\mathcal{O}_{C}^{\rm{(SM)}}=\overline{L_{e}}He_{R}$ and $\mathcal{O}_{X}^{(D)}=X^{p}$ in different temperature regimes. The first column shows $T_{\rm{dec}}$, which is identified with $T_{B-L}$ $(T_{\rm{eq}})$ for $\bar{\mu}_{B-L}^{\rm eq}$ $\left(\mu_{X}^{\rm eq}\right)$, see text below Eq. (6), and (i)-(v) indicate the following ranges for $T_{\rm{dec}}$ [GeV]: $\quantity(10^{5},10^{6})$, $\quantity(10^{6},10^{9})$, $\quantity(10^{9},10^{11-12})$, $\quantity(10^{11-12},10^{13})$, $\quantity(10^{13},10^{15})$, respectively. The flavor index $\alpha$ in the second column lists the uncorrelated charged-lepton flavors that interact with $N_{1}$ in the relevant temperature range, assuming the SM–DM portal is active (decoupled) in that range. In the latter case, the results for $x_{C}$ reduce to those in standard WILG, c.f. Table II in Domcke et al. (2021), which are included in parentheses in column 33 below. The ✗ symbol marks the absence of the corresponding $\bar{\mu}_{C}$ due to an efficient SM interaction. The coefficients $P$ and $P_{e,\tau}$ encode the primordial $q_{e,\mu,\tau}$ asymmetries’ flavor composition relative to the $N_{1}$ washout direction, as detailed in Domcke et al. (2021). See text for further details.
$T_{\rm{dec}}$ Index $\alpha$ $\bar{\mu}_{Xe}$ ($\bar{\mu}_{e}$) $\bar{\mu}_{2B_{1}-B_{2}-B_{3}}$ $\bar{\mu}_{u-d}$ $\bar{\mu}_{d-s}$ $\bar{\mu}_{B_{1}-B_{2}}$ $\bar{\mu}_{\mu}$ $\bar{\mu}_{u-c}$ $\bar{\mu}_{\tau}$ $\bar{\mu}_{d-b}$ $\bar{\mu}_{B}$ $\bar{\mu}_{u}$ $\bar{\mu}_{\Delta_{\perp}}$ $\bar{\mu}_{\Delta_{\alpha}}$
(i) $e,\mu,\tau$ $(e,\mu,\tau)$ $\frac{-3p^{2}}{22+10p^{2}}$ $(\frac{-3}{10})$$\frac{-711}{1422+481p^{2}}$ ✗ (✗), ✗ (✗), ✗ (✗), ✗ (✗), ✗ (✗), ✗ (✗), ✗ (✗), ✗ (✗), ✗ (✗), ✗ (✗), ✗ (✗), $\frac{[-185,\,52,\,52]}{1422+481p^{2}}$
(ii) $e,\mu,\tau$ $(e,\mu,\tau)$ $\frac{-3p^{2}}{38+17p^{2}}$ $\left(\frac{-3}{17}\right)$$\frac{-1071}{2142+716p^{2}}$ 0 $(0)$, 0 $\frac{-16-7p^{2}}{38+17p^{2}}$ $\left(\frac{-7}{17}\right)$, $\frac{45}{2142+716p^{2}}$ ✗ (✗), ✗ (✗), ✗ (✗), ✗ (✗), ✗ (✗), ✗ (✗), ✗ (✗), ✗ (✗), ✗ (✗), $\frac{[-265,\,92,\,92]}{2142+716p^{2}}$
(iii) $e,\mu,\tau$ $(\parallel_{\tau},\tau)$ $\frac{-3p^{2}}{62+27p^{2}}$ $\left(\frac{142-225P_{\tau}}{247}\right)$$\frac{-1779}{3558+1178p^{2}}$ 0 (0)(0),0 $\frac{-6(7+3p^{2})}{62+27p^{2}}$ $\left(\frac{-123}{247}\right)$$\frac{135}{3558+1178p^{2}}$ $\frac{-4(7+3p^{2})}{62+27p^{2}}$$\left(\frac{-82}{247}\right)$$\frac{45}{1779+589p^{2}}$ $\frac{21+9p^{2}}{62+27p^{2}}$ $\left(\frac{123}{494}\right)$$\frac{-135}{7116+2356p^{2}}$ $\frac{-3(2+p^{2})}{62+27p^{2}}$ $\left(\frac{142-225P_{\tau}}{247}\right)$$\frac{-6}{1779+589p^{2}}$ ✗ (✗), ✗(✗), ✗(✗), ✗(✗), ✗(✗), $\left(\frac{225}{247}\right)$, $\frac{[-421,\,168,\,172]}{3558+1178p^{2}}$
(iv) $e,\parallel_{e}$ $(\parallel)$ $\frac{-9p^{2}}{102+37p^{2}}$ $\left(\frac{-23P+7}{30}\right)$$\frac{-38}{76+23p^{2}}$ $\frac{12(3+p^{2})}{102+37p^{2}}$ $\left(\frac{1}{5}\right)$$\frac{-3}{76+23p^{2}}$ $\frac{-36(3+p^{2})}{102+37p^{2}}$ $\left(\frac{-3}{5}\right)$$\frac{9}{76+23p^{2}}$ $\frac{-10(3+p^{2})}{102+37p^{2}}$ $\left(\frac{-1}{6}\right)$$\frac{5}{152+46p^{2}}$ $\frac{-18(3+p^{2})}{102+37p^{2}}$ $\left(\frac{-3}{10}\right)$$\frac{9}{152+46p^{2}}$ $\frac{42+p^{2}(14-23P_{e})-60P_{e}}{102+37p^{2}}$ $\left(\frac{-23P+7}{30}\right)$$\frac{-7}{152+46p^{2}}$ $\frac{18(3+p^{2})}{102+37p^{2}}$ $\left(\frac{3}{10}\right)$$\frac{-9}{152+46p^{2}}$ $\frac{42+p^{2}(14-23P_{e})-60P_{e}}{102+37p^{2}}$ $\left(\frac{-23P+7}{30}\right)$$\frac{-7}{152+46p^{2}}$ $\frac{-16(3+p^{2})}{102+37p^{2}}$ $\left(\frac{-4}{15}\right)$$\frac{4}{76+23p^{2}}$ $\frac{120+46p^{2}}{306+111p^{2}}$ $\left(\frac{23}{90}\right)$$\frac{23/6}{76+23p^{2}}$ ✗ (✗), $\frac{60+23p^{2}}{102+37p^{2}}$ $\left(\frac{23}{30}\right)$$\frac{7}{152+46p^{2}}$ $\frac{[-16,7]}{152+46p^{2}}$
(v) $e,\parallel_{e}$ $(\parallel)$ $\frac{-p^{2}}{14+5p^{2}}$ $\left(\frac{-3P+1}{4}\right)$$\frac{-5}{10+3p^{2}}$ $\frac{4(3+p^{2})}{42+15p^{2}}$ $\left(\frac{1}{6}\right)$, $\frac{-1}{30+9p^{2}}$ $\frac{-20(3+p^{2})}{42+15p^{2}}$$\left(\frac{-5}{6}\right)$$\frac{5}{30+9p^{2}}$ $\frac{-2(3+p^{2})}{14+5p^{2}}$ $\left(\frac{-1}{4}\right)$, $\frac{1}{20+6p^{2}}$ $\frac{-2(3+p^{2})}{14+5p^{2}}$ $\left(\frac{-1}{4}\right)$$\frac{1}{20+6p^{2}}$ $\frac{6+p^{2}(2-3P_{e})-8P_{e}}{14+5p^{2}}$ $\left(\frac{-3P+1}{4}\right)$$\frac{-1}{20+6p^{2}}$ $\frac{2(3+p^{2})}{14+5p^{2}}$ $\left(\frac{1}{4}\right)$$\frac{-1}{20+6p^{2}}$ $\frac{6+p^{2}(2-3P_{e})-8P_{e}}{14+5p^{2}}$ $\left(\frac{-3P+1}{4}\right)$, $\frac{-1}{20+6p^{2}}$ $\frac{-8(3+p^{2})}{42+15p^{2}}$ $\left(\frac{-1}{3}\right)$$\frac{2}{30+9p^{2}}$ $\frac{10+4p^{2}}{42+15p^{2}}$ $\left(\frac{1}{6}\right)$$\frac{2}{30+9p^{2}}$ $\frac{8(3+p^{2})}{42+15p^{2}}$ $\left(\frac{1}{3}\right)$, $\frac{-2}{30+9p^{2}}$ $\frac{8+3p^{2}}{14+5p^{2}}$ $\left(\frac{3}{4}\right)$$\frac{1}{20+6p^{2}}$ $\frac{[-2,\,1]}{20+6p^{2}}$
@@
Asymgenesis —  Consider a general scenario where primordial charges in both the visible and the dark sector can take arbitrary values and the charge-transfer interaction is in thermal equilibrium at high temperatures. Schematically, this corresponds to the regime above the barrier in the right panel of Fig. 1. The BLB-L asymmetry in the strong wash-in regime and the ADM asymmetry can then, in full generality, be expressed as
@@
where $s$ denotes entropy density, and $T_{B-L}$ $(T_{\rm{eq}})$ denotes the temperature when RHN interactions (the charge-transfer interaction) become(s) inefficient. Here, $\Delta_{\alpha}=B/3-L_{\alpha}$ denote lepton-flavor asymmetries corresponding to the flavor combinations that are decohered at the relevant temperature scale. At temperatures below $T_{\rm{eq}}$, corresponding to the regime below the top of the barrier in Fig. 1, both $q_{e}$ and $q_{X}$ become separately conserved. In particular, for $T_{B-L}<T_{\rm{eq}}$, $C=Xe$ can be replaced with $C=e$ in Eq. (6), with $q_{e}$ given by
@@
In this case, the numerical values for $x_{C}$ reduce to the result from standard WILG Domcke et al. (2021). Figure 2 provides an overview of the thermal history in Asymgenesis, highlighting the characteristic temperature regimes and the evolution of charge asymmetries. The left panel shows the scenario where $T_{B-L}>T_{\rm{eq}}$, while the right panel shows the opposite ordering $T_{\rm{eq}}>T_{B-L}$. Moreover, the values for $x_{C}$ and $y_{C}$ are shown for various temperature regimes in Table 1 for our example where $\mathcal{O}_{C}^{\rm{(SM)}}=\overline{L_{e}}He_{R}$ [^4] It is worth noting that if the second or third generation of leptons should enter $\mathcal{O}_{C}^{\rm{(SM)}}$ instead of the first generation as considered here, then the charge-transfer interaction must decouple above the equilibration temperatures of the $\mu$ or the $\tau$ Yukawa interactions $T_{\mu}=10^{9}$ GeV or $T_{\tau}=10^{12}$ GeV, respectively, to avoid large washout of the asymmetric DM component. A detailed account of asymmetry generation for different choices of $\mathcal{O}_{C}^{\rm{(SM)}}$ will be presented elsewhere ”Mojahed, Martin A. and Weber, Sascha: Work in progress” .. The coefficients without parentheses correspond to $T_{\rm{CG}}>T_{B-L}>T_{\rm{eq}}$, while the coefficients in parentheses correspond to $T_{\rm{CG}}>T_{\rm{eq}}>T_{B-L}$, where $T_{\rm{CG}}$ is the temperature scale associated with CG. The table shows that both the asymmetry in $X$ particles and the BAU are typically related to the primordial charge asymmetries generated during CG by an $\mathcal{O}(0.1-1)$ number. The analogous statement in the context of standard ADM models is that if the BLB-L and $X$-violating higher-dimensional operator is in thermal equilibrium at high temperatures, then the BAU is related to the asymmetric component of DM by a $\mathcal{O}(1)$ number, barring potential washout processes. In conclusion, the framework considered so far predicts DM masses of $\mathcal{O}(1)-\mathcal{O}(10)$ GeV, which is similar to standard ADM models.
Beyond thermal equilibrium: SM to DM — Up to this point, we have analyzed the exchange of charges between the SM and the dark sector under the assumption that the charge-transfer interaction attained thermal equilibrium in the early universe. This requirement, however, is not essential. The same dynamics can be treated in an out-of-equilibrium framework using semiclassical Boltzmann equations. In this section, we relax the assumption of thermal equilibrium and consider a scenario in which the primordial charges in the dark sector are negligible. In the next section, we study the opposite limit, where the primordial charges in the SM sector are negligible. More generally, however, we emphasize that a CG mechanism can generate nonzero primordial charges in both the SM and the dark sector. To facilitate this description, we introduce the following quantities
@@
The washout term, $W(z)$, associated with the interactions induced by Eq. (1) takes the following form, [^5] The expression for $\Gamma^{w}$ is only valid in the regime where $T<\Lambda$, while for $T\gtrsim\Lambda$ it would be expected to take the following form $\Gamma^{w}=cT$, where $c$ is a constant.
@@
[…]
@@
[…]
and $z_{\rm{CG}}=M_{N}/T_{\rm{CG}}$. As the electron-Yukawa interaction comes into equilibrium, $Y_{\Delta X}^{\rm{eq}}$ goes to zero. If, at that stage, the charge-transfer interaction characterized by $\Gamma^{w}$ remains efficient, Eqs. (12) and (13) show that $Y_{\Delta X}$ is driven to zero, thereby washing out the $X$ asymmetry. Meanwhile, if the charge-transfer interactions are initially in thermal equilibrium and decouple above the electron Yukawa equilibration temperature, $T_{e}\approx 85$ TeV Bödeker and Schröder (2019), then the final ADM yield, $Y_{\Delta X}^{\rm{end}}$, is driven exponentially close to $Y_{\Delta X}^{\rm{eq}}$, as given by Eq. (10) and Table 1. The attraction of $Y_{\Delta X}^{\rm{end}}$ towards $Y_{\Delta X}^{\rm{eq}}$ driven by $W$, can be referred to as dark wash-in, see also Asadi et al. (2025).
Suppose instead that the effective-operator scale is so high that the charge-transfer interaction never enters thermal equilibrium. The charge transfer then becomes a UV freeze-in (FI) process Giudice et al. (2001); Hall et al. (2010); Elahi et al. (2015). In such a case, $Y_{\Delta X}^{\rm{end}}$ is to a good approximation, determined by the value of $Y_{\Delta X}^{\rm{eq}}$ right after CG, and the value of $\Lambda$ and $T_{\rm{CG}}$. To unify the UV FI regime and the thermal-equilibrium regime, we find it useful to express the final asymmetry in $\Delta X$ as
@@
where $z^{\prime}=\text{Max}[z_{\rm{CG}},z_{\text{eq}}]$ and $\kappa$ parametrizes the charge-transfer efficiency. Numerical values for $\kappa$ for the case, $\mathcal{O}_{X}^{(D)}=X^{2}$, are shown in Fig. 3, where the horizontal and vertical axes show $T_{\rm{CG}}$ and $\Lambda$, respectively, and the black solid curve shows $H(z_{\rm{CG}})=\Gamma^{w}(z_{\rm{CG}})$. The latter marks the transition between the two regimes: The area below the curve corresponds to the thermal-equilibrium regime, $\Gamma^{w}(T_{\rm CG})/H(T_{\rm CG})>1$, where $\kappa\simeq 1$ and the asymmetry is well described by Tab. 1. The area above the curve shows the UV FI regime, $\Gamma^{w}(T_{\rm CG})/H(T_{\rm CG})<1$, and in the limit $\Gamma^{w}(T_{\rm CG})/H(T_{\rm CG})\ll 1$, we have $\kappa\simeq\Gamma^{w}(T_{\rm CG})/H(T_{\rm CG})$. For charge transfer from the SM to the DM sector, the asymmetry in the dark sector is significantly smaller than the BAU in large portions of parameter space where Asymgenesis proceeds via UV FI. The black-dashed curve shows the contour corresponding to $\kappa=10^{-3}$, which would point to DM masses of $\mathcal{O}(1)$ TeV. This roughly represents an upper bound on ADM masses when the symmetric annihilation cross section is required to be large enough for the ADM component to be the dominant contribution to the DM relic density while respecting perturbative unitarity Griest and Kamionkowski (1990); Cohen et al. (2010); Petraki and Volkas (2013). Interestingly, for $T_{\rm{CG}}=10^{14-15}$ GeV, it could be possible to account for the DM abundance even if the scale of NP associated with charge transfer is related to the scale of grand unification or possibly quantum-gravity effects, c.f. the upper-right corner of the figure. In summary, the region below the black solid curve in Fig. 3 corresponds to DM masses in the range of $\mathcal{O}(1)-\mathcal{O}(10)$ GeV, while the area below the black dashed curve corresponds to DM masses up to $\mathcal{O}(1)$ TeV. Above the black dashed curve, the symmetric component cannot annihilate efficiently due to the constraints imposed by perturbative unitarity. Finally, if the symmetric component cannot be annihilated efficiently, the dark matter mass must be smaller than the values quoted here to avoid overclosing the universe.
[…]
@@
[…]
@@
[…]
@@
[…]
Finally, we note that since Asymgenesis builds on the type-I seesaw extension of the SM, it might be interesting to apply the framework to BSM models with gauged $U(1)_{B-L}$ symmetry. Contrary to conventional ADM models, Asymgenesis can accommodate scenarios where the charge-transfer mechanism decouples at temperatures that are orders of magnitude above the BLB-L-breaking scale, and still result in comparable abundances of baryonic matter and ADM.
Conclusion — In this work, we have proposed Asymgenesis, a framework based on the type-I seesaw model that can be used to account for the BAU and the observed DM density from initial conditions appropriate to either ADM or WILG. Compared to standard ADM scenarios that rely on higher-dimensional operators, Asymgenesis offers significantly greater flexibility in the charge-transfer interactions between the dark and visible sectors. Crucially, it disentangles the temperature at which the charge-transfer interactions decouple from the temperature scale at which BLB-L violation becomes effective. By combining the advantages of WILG and ADM, the Asymgenesis framework considerably widens the model-building landscape for asymmetric dark matter.